Two dices are thrown simultaneously to get the co-ordinates of a point on(x,y) plane. Find the probabilty taht this point lies on or inside the region bounded by |x|+|y|=3 .

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Debjit,

{x,y}={1,2},{2,1},{1,1}

answer=2*(1/62)+(1/62)

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Debjit !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR

Ashwin Muralidharan IIT Madras
12 years ago

Hi Debjit,

 

Out of the total possibilities of 36 outcomes, the ones that satisfy |x|+|y| = 3,

would be (1,1) (1,2) (2,1).

 

So 3 favourable ways out of a total of 36 outcomes.

 

Hence Probability = 3/36 = 1/12.

 

Hope this helps.

 

Best Regards,

Ashwin (IIT Madras).